find the standard form of the equation of a circle with endpoints of a diameter at the points (7,2) (-9,5)
Answer by nshah11(47)

Center is midpoint of endpoints => ((7 - 9)/2, (2 + 5)/2)= (-1, 7/2)
Radius^2 = 64 + 9/4 = 265/4
So equation of circle is:
(x + 1)^2 + (y - 7/2)^2 = 265/4